OpenStudy (anonymous):

If the hot dog eating champion's goal is to eat 1 hot dog every 13 seconds, how many whole hot dogs does he expect to eat in 12 minutes? A. 156 hot dogs B. 55 hot dogs C. 56 hot dogs D. 48 hot dogs

OpenStudy (anonymous):

a i think

OpenStudy (anonymous):

12*60=720 720/13=?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@katty420

OpenStudy (anonymous):

No, 720/13=55.3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

A=156, and the answer is 55.3. Round 55.3 to the nearest whole number, and you won't get a.

OpenStudy (anonymous):

thx